OpenStudy (anonymous):

The profit made from the sale of tiles is found by subtracting the costs from the revenue. e.Find the Profit Equation by substituting your equations for R and C in the equation . Simplify the equation. f.What is the profit made from selling 20 tile sets per month? g.What is the profit made from selling 25 tile sets each month? h.What is the profit made from selling no tile sets each month? Interpret your answer. i.Use trial and error to find the quantity of tile sets per month that yields the highest profit. j.How much profit would you earn from the number you found in part i? k.Wh

OpenStudy (anonymous):

k. What price would you sell the tile sets at to realize this profit? Hint: Use the demand equation from part a.

OpenStudy (anonymous):

\[\Pi = R - C\] Just plug in the equations you got for R and C

OpenStudy (anonymous):

im still confused.

OpenStudy (anonymous):

\[R = -x + 62\]\[C = 6x + 300\]\[\Pi = (-x + 62) - (6x + 300) = -x + 62 - 6x - 300\]